Grade 11Physical Chemistry

Calculate molar energy in kj corresponding to photons of energy 2 ev?

Rishabh Singh
10 Years agoGrade 11

2 Answers

Ashutosh Mohan Sharma
10 Years ago
energy of photon = 2 ev
energy of 1 mole of photons = 2 x 6.022 x 10^23 = 1.02044 x 10^24 ev
= 1.02044 k10^24 x 1.6 x 10^-19
= 1.6327 x 10^5 J
= 1632.7 KJ
Akshay
8 Years ago
Energy of photon = 2 eV 1 mole photon = 6.022 x 10^23 no. of photonstherefore, molar energy of given photons = 2 x 6.022 x 10^23 eV = 12.044 x 10^23 eV = 12.044 x 10^23 x 1.6 x 10^-19 J (Since, 1 eV = 1.6 x 10^-19) = 19.2704 x 10^4 J = 192.704 KJ (Since, 10^3 J = 1 KJ) = 193 KJ (approximate molar energy)